OpenStudy (barrelracer011):

If the area of a rectangle 24a2b and the length is 8ab2, what would the width of the rectangle, given that width is found by dividing area by length? Simplify the answer. the quantity of 3 times a all over b b over the quantity 3 times a 3 over the quantity a times b 2st

OpenStudy (anonymous):

Divide: 24(a^2)b / 8a(b^2) = 24/8 * a^2/a * b/b^2 = ?

OpenStudy (anonymous):

is it \(A=24a^2b,\ and\ l=8ab^2\)? if so, \(\LARGE w=\frac{A}{l}=\frac{24a^2b}{8ab^2}\) remember the law of exponents \(\Large \frac{x^m}{x^n}=x^{m-n}\)

OpenStudy (anonymous):

24a2b8ab2=248∗a2a∗bb2=?

OpenStudy (barrelracer011):

3?

OpenStudy (anonymous):

you cannot eliminate the variables

OpenStudy (barrelracer011):

I know Iam redoing it to figure out where the variables go....

OpenStudy (barrelracer011):

Ok so it would be A

OpenStudy (barrelracer011):

@Chibi_Robo3

OpenStudy (anonymous):

yep ;)

OpenStudy (barrelracer011):



OpenStudy (barrelracer011):

Thanks!